OpenStudy (xmoses1):

Which of the following represents the general term for the sequence 2, 4, 6, 8, 10, . . .? n + 1 2n 2n - 1

OpenStudy (xmoses1):

@AravindG This is it!

OpenStudy (anonymous):

well do it like this! n starts at 1 right? n+1 = 2 @ n=1 good. n+1 = 4 @n=2 no good.

OpenStudy (aravindg):

First identify what type of sequence it is.

OpenStudy (xmoses1):

It seems to be Arithmetic...

OpenStudy (aravindg):

Good. Do you know general term for an Arithmetic sequence?

OpenStudy (anonymous):

2n = 2 @ n=1 good. what about the rest of these numbers?

OpenStudy (xmoses1):

Not off the top of head. I think it is something like \( {a_n} {(2^n +1)} =1 \) right? Or close to that

OpenStudy (aravindg):

\[a_n=a+(n-1)d\] a=first term d=common diff

OpenStudy (aravindg):

Substitute and be happy :)

OpenStudy (xmoses1):

Ah I see! Thanks :) Let me see... Then it would be 2n right?

OpenStudy (aravindg):

yes :)

OpenStudy (xmoses1):

Thanks @AravindG :D

OpenStudy (aravindg):

yw :) But this is the general method. The answer could have been just got by looking at options too. Just applying common sense, 2n represents 2,4,6,8 etc.

OpenStudy (xmoses1):

Yeah that is what i was having issues with lol. On the other ones i could do that, but this one i was getting very well. I am also going to choose not to take that as an insult to my common sense level xD

OpenStudy (xmoses1):

*Was not